Grade 12th passAlgebra

Find the area of the following circles, given that:
Radius = 5 cm

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given, radius of circle = 5 cm
Then,
Area of the circle =πr2
= 22/7 × 52
= 22/7 × 25
= 550/7
= 78.57 cm2

Thanks